kelly is 1/3 the age of her father. in 12 years she will be half the age of her father. how old is each now?
Answer by jorel1380(3719)

f=3k
f+12=2(k+12)
3k+12=2k+24
k=12
f=36
Kelly is 12. Her father is 36..